How many moles of Sn4+ are reduced to Sn2+ by 1.00 C of electric charge?


1
Expert's answer
2021-02-15T03:18:55-0500

1) 1.00C = 6.24 * 10^18 electrons.

2) due to the fact that each atom was reduced by 2 electrons, the quantity of atoms equals:

N = 1/2 * 6.24 * 10^18 = 3.12 * 10^18

3)"n = \frac{N}{N\tiny{A}};" n = 3.12 * 10^18 / 6.022 * 10^23 = 0.518 * 10^(-5) = 5.18 * 10^(-6)